# Certifying rings of integers in number fields

### v1

@@ 4.7. Global maximality
Let $K$ be the number field obtained by adjoining a root $\theta$ of the monic polynomial $T$ in $\mathbf{Z}[X]$. We aim to verify that a subalgebra $\mathcal{O}$, with an explicitly given basis, is equal to $\mathcal{O}_{K}$.
Since $T$ is separable, we can find $a,b\in\mathbf{Z}[X]$ such that $aT+bT^{\prime}=n$ for some nonzero integer $n$. If $p$ does not divide $n$, then $T\pmod{p}$ is squarefree and it easily follows that $T$ satisfies Dedekind criterion at $p$. Given $\theta\in\mathcal{O}$, this gives a quick way to prove $p$-maximality of $\mathcal{O}$ for all but finitely many primes $p$ (those dividing $n$). In fact, $T$ may still satisfy the criterion for some of these primes. For the rest, the certificates in Section 4.6 can be used, leading to a proof of $\mathcal{O}=\mathcal{O}_{K}$.
@@
where $K$ is AdjoinRoot (map (algebraMap ℚ) T), and $\mathcal{O}$ is the subalgebra of $K$ with integral basis $\{w_{i}\}_{i}$, constructed as in Section 4.1. The proof of irreducibility of $T$ is generated as in Section 3. Lean can then infer a NumberField K instance. The term NumberField.RingOfIntegers K is simply the coercion of IntegralClosure K into a type.
@@ 5. Resultants and discriminants
Consider an $R$-algebra $A$ with a finite $R$-basis. Given an element $x\in A$, the map $y\mapsto xy$ is an $R$-linear map, and thus can be represented as a matrix with entries in $R$. The trace of this matrix, which is independent of the choice of basis, is known as the trace of $x$ and denoted by $\operatorname{Tr}_{A/R}(x)$. For a collection of elements $a_{1},\ldots,a_{n}$ of $A$, its discriminant is defined by $\operatorname{disc}(a_{1},\ldots,a_{n})=\det(\operatorname{Tr}_{A/R}(a_{i}a_{j}))_{ij}$. When we started our project, Mathlib already contained this definition.
If $K=\mathbf{Q}(\alpha)$ is a number field of degree $n$ and $\mathcal{O}$ is a subring which is free of rank $n$ as a $\mathbf{Z}$-module, we define $\operatorname{disc}(\mathcal{O})\in\mathbf{Z}$ as the discriminant of an integral basis for $\mathcal{O}$. In particular, $\operatorname{disc}(\mathcal{O}_{K})$ is known as the discriminant of the number field $K$ and it is an important arithmetic invariant encoding information about $K$. If $\alpha\in\mathcal{O}$, then we have
@@
The index $[\mathcal{O}:\mathbf{Z}[\alpha]]$ can be determined from an explicit basis for $\mathcal{O}$ in terms of $\alpha$. Knowing the discriminant of the power basis $1,\alpha,\ldots,\alpha^{n-1}$ then helps us compute $\operatorname{disc(\mathcal{O})}$.
@@
In the remainder of this section, let $R$ be an integral domain, $f,g$ be polynomials of degree $n$ and $m$ respectively with coefficients given by $f(X)=a_{n}X^{n}+\cdots+a_{1}X+a_{0}\in R[X]$ and $g(X)=b_{m}X^{m}+\cdots+b_{1}X+b_{0}\in R[X]$. Our path towards formalized computation of the discriminant starts with the notion of the resultant $\operatorname{Res}(f,g)$ defined as the determinant of the Sylvester matrix of dimension $(m+n)\times(m+n)$, determined by the coefficients of $f$ (see (Weisstein, 2024) for a transposed version with $f$ and $g$ interchanged). We define the discriminant of a polynomial $f$ as $\operatorname{Disc}(f)=\operatorname{Res}(f,f^{\prime})$ where $f^{\prime}$ is the formal derivative of $f$. This differs slightly from the usual definition, which includes a scaling factor depending on the leading coefficient and the degree of $f$.
@@
$\operatorname{Res}(f,g)=(-1)^{nm}a_{n}^{m}b_{m}^{n}\prod_{i}\prod_{j}(\alpha_{i}-\beta_{j}).$
Note that the assumption of Theorem 5.1 always applies after mapping $f$ and $g$ to the algebraic closure of the fraction field of $R$, and this map will preserve the resultant, so we will assume that $f$ and $g$ split completely. It then follows from the product rule and elementary algebraic manipulations that $\prod_{i}\prod_{j>i}(\alpha_{j}-\alpha_{i})^{2}=(-1)^{n(n-1)/2}\operatorname{Disc}(f)$. Since the former term is shown in Mathlib to equal the discriminant of the power basis $1,\ldots,\alpha^{n-1}$, and the latter term is computable, this gives us a way to compute, given a basis for $\mathcal{O}_{K}$ and using (5), the discriminant of a number field.
@@
Next, we viewed $\operatorname{Res}(f,g)$ and $\prod_{i}\prod_{j}(\alpha_{i}-\beta_{j})$ as multivariate polynomials in $R[\alpha_{1},\dots,\alpha_{n},\beta_{1},\dots,\beta_{m}]$. We showed that these are indeed homogeneous polynomials, of degree $mn$, where an essential ingredient was characterizing the homogeneous polynomials $p\in R[x,y,\dots,z]$ as those polynomials where $p(cx,cy,\dots,cz)=c^{k}p(x,y,\dots,z)$, assuming $R$ is infinite. The coefficient of $(\alpha_{1}\alpha_{2}\cdots\alpha_{n})^{m}$ in $\operatorname{Res}(f,g)$ and in $\prod_{i}\prod_{j}(\alpha_{i}-\beta_{j})$ equals $1$, so if the two expressions are equal up to multiplication by $c$, we must have $c=1$ and the expressions are actually equal.
@@
Although this approach does not differ significantly from one found in textbooks such as Lang’s Algebra (Lang, 2002), a particular source of difficulty we had to deal with was the textbooks’ implicitly shifting point of view between different polynomial rings. For example, Lang argues that $\operatorname{Res}(f,g)$ is homogeneous of degree $m$ as a polynomial in $R[a_{0},\dots,a_{n}]$ and of degree $n$ as a polynomial in $R[b_{0},\dots,b_{m}]$, but that $\alpha_{i}-\beta_{j}$ divides $\operatorname{Res}(f,g)$ in $R[\alpha_{1},\dots,\alpha_{n},\beta_{1},\dots,\beta_{m}]$. Our formalization works exclusively in the polynomials
$R[\alpha_{1},\dots,\alpha_{n},\beta_{1},\dots,\beta_{m}]$.
To evaluate discriminants, we use the Polynomial.ofList definition of Section 3.1. Expressions such as resultant (ofList [5, 4, 3, 2, 1]) (ofList [4, 6, 6, 4]) are computable, although we run into a subsequent error: the implementation of determinants in Mathlib is not optimized for computation and therefore causes a stack overflow. However, we were still able to compute the discriminant of degree $3$ polynomials such as $f=X^{3}-30X-80$ and, after formally verifying an integral basis for the ring of integers of $K=\mathbf{Q}[X]/\langle f\rangle$, proved that
@@ 6. Formally verifiying LMFDB entries
A popular publicly available online database containing number theoretic data is the L-functions and modular forms database (LMFDB) (LMFDB Collaboration, 2024). It gathers data from various sources, relying on several programming languages and CASs. Much of the data is currently beyond formally verifiable, but our work makes it possible to formally verify some nontrivial entries for number fields of arbitrary degree. In particular, we focus on checking the integral basis for a choice of number fields up to degree $8$. To a lesser extent, we also compute the corresponding discrimiminant, but only up to degree $3$, as our computational methods are currently still under development. The aim at this point is not to formally verify integral bases and discriminants for a large fraction of the LMFDB, but to show the feasibility, and current computational limits, of the approach.
Concretely, there are $142$ number fields in the LMFDB which are of degree $5$, unramified outside $2,3,5$, and have (to focus on complications) no monogenic ring of integers. (This list, up to isomorphism, is actually complete, but this is not our concern here.) For all these $142$ number fields we successfully used our tools to formally verify the given integral basis. The most time consuming step when checking these proofs appears to be the exponentiation step in the certificates of Section 4.6. We recall that this is done with a recursively defined function which does repeated squaring and multiplication, using the times table of $\mathcal{O}/p\mathcal{O}$. In three cases, when the prime involved exceeded 400, checking the proof, which consisted of repeated rfl, initially failed due to reaching the default number of maximum heartbeats of 200,000. Manually unfolding the definition of the exponentiation function a couple of times before repeating rfl allowed the proofs to typecheck without increasing the heartbeat limit.
As our discriminant computations cannot yet handle large degree, we also tested degree $3$ cases. There are $7$ number fields in the LMFDB which are of degree $3$, unramified outside $2,3,5$, and have no monogenic ring of integers. For all these number fields we successfully formally verify the given integral basis as well as the discriminant. We note that in all of these cases, the longer form of the certificate for $p$-maximality (given in detail in Appendix A) had to be used. In contrast, for all of degree 5 examples except one, the shorter certificate with a witness was sufficient.
We also successfully verified a degree 6 example and a simpler, degree 8 example, increasing the heartbeats for the latter. In this last example, the system struggled finding the instance AddCommMonoid for the subalgebras $\mathcal{O}$ and $\mathcal{O}_{m}$, requiring additional typeclass synthesis heartbeats.
@@ 7.1. Future work
The discriminant computations should be improved. At this point, these are directly computed from the corresponding resultant. For repeated use of discriminant evaluation of a fixed degree, general formulae for the discriminant of a degree $n$ polynomial should be formalized, up to some ‘reasonable’ $n$. The computations for the ring of integers are in a more advanced stage, though further optimization is needed to deal with higher degrees in order to be able to cover all number fields from the LMFDB. We note that much of the mathematical foundation is formalized in enough generality to allow for possible extension to more general contexts. Finally, certifying other fundamental invariants, such as class groups and unit groups, would be very interesting. While these would constitute completely new projects, they could build on the formalized rings of integers.
@@ 7.2. Related work
As mentioned above, some rings of integers for quadratic number fields have been formalized before (Baanen et al., 2023); this concerns both the determination as well as concrete computation. Specific instances, such as the Gaussian integers $\mathbf{Z}[i]$, have been considered in various systems; see e.g. (Eberl, 2020; Futa et al., 2012). Mathematical certificates outside of algebraic number theory have received much attention. Some examples relevant for (general) number theory and computer algebra include (Thiemann et al., 2020; Mahboubi and Sibut-Pinote, 2021). Previous work bridging Lean with computer algebra systems includes the tactics polyrith, written by Dhruv Bhatia, and sageify (Baanen et al., 2023), both of which integrate with SageMath. Additionally,  (Lewis, 2017; Lewis and Wu, 2022) develop a connection between Lean and Mathematica.
@@ 7.3. Conclusions
In total, for our developments, we wrote about 13000 lines of Lean code. Our SageMath scripts generated about 32000 lines of Lean code for the certifications we considered. With these kind of projects, it is very hard to reliably measure the De Bruijn factor.
Appendix A Appendix
The following certificate is guaranteed to always exist if $K$ is a number field, $\mathcal{O}_{m}=\mathcal{O}_{K}$, and $\mathcal{O}$ is $p$-maximal.
@@ Appendix A Appendix
Consider $\mathcal{O}$ and $\mathcal{O}_{m}$ subalgebras of a $\mathbf{Z}$-torsion free commutative ring such that $\mathcal{O}\subseteq\mathcal{O}_{m}$, and both of rank $r$. Let =$\mathcal{B}=\{b_{1},\ldots,b_{r}\}$ be a $\mathbf{Z}$-basis for $\mathcal{O}$. Let $m,n$ and $t$ be non-negative integers with $r=m+n$ and $r\leq p^{t}$. A certificate for $p$-maximality consists of the following data:
@@
(p : ℕ) [hpI : Fact \$ Nat.Prime p] ($\mathcal{O}$ : Subalgebra K) ($\mathcal{O}$m : Subalgebra K) (hm : $\mathcal{O}$ $\mathcal{O}$m) where

### v2

@@ 5.7. Global Maximality
Here, we describe how we can automatically generate a Lean proof of $\mathcal{O}=\mathcal{O}_{K}$ using the certificates previously introduced.
Let $K$ be the number field obtained by adjoining a root $\theta$ of the monic irreducible polynomial $T$ in $\mathbf{Z}[X]$. We aim to verify that a subalgebra $\mathcal{O}$, with an explicitly given basis, is equal to $\mathcal{O}_{K}$. Since $T$ is separable, we can find $a,b\in\mathbf{Z}[X]$ such that $aT+bT^{\prime}=n$ for some nonzero integer $n$. If $p$ does not divide $n$, then $T\pmod{p}$ is squarefree and it easily follows that $T$ satisfies the Dedekind criterion at $p$. Given $\theta\in\mathcal{O}$, this gives a quick way to prove $p$-maximality of $\mathcal{O}$ for all but finitely many primes $p$ (those dividing $n$). In fact, $T$ may still satisfy the criterion for some of these primes. For the rest, the certificates in Section 5.6 can be used, leading to a proof of $\mathcal{O}=\mathcal{O}_{K}$.
@@
where $K$ is AdjoinRoot (map (algebraMap ℚ) T), and $\mathcal{O}$ is the subalgebra of $K$ with integral basis $\{w_{i}\}_{i}$, constructed as in Section 5.1. The proof of irreducibility of $T$ is generated as in Section 4. Lean can then infer a NumberField K instance. The term NumberField.RingOfIntegers K is simply the coercion of IntegralClosure K into a type. As mentioned, while generating Lean code from SageMath was sufficient for our purposes, tighter integration between the CAS and Lean remains an interesting direction for future work.
@@ 6. Resultants and Discriminants
The goal of this section is to provide a method for computing the discriminant of an integral basis of $\mathcal{O}$. Consider an $R$-algebra $A$ with a finite $R$-basis. Given an element $x\in A$, the map $y\mapsto xy$ is an $R$-linear map, and thus can be represented as a matrix with entries in $R$. The trace of this matrix, which is independent of the choice of basis, is known as the trace of $x$ and denoted by $\operatorname{Tr}_{A/R}(x)$. For a collection of elements $a_{1},\ldots,a_{n}$ of $A$, its discriminant is defined by $\operatorname{disc}(a_{1},\ldots,a_{n})=\det(\operatorname{Tr}_{A/R}(a_{i}a_{j}))_{ij}$. When we started our project, Mathlib already contained this definition.
If $K=\mathbf{Q}(\alpha)$ is a number field of degree $n$ and $\mathcal{O}$ is a subring which is free of rank $n$ as a $\mathbf{Z}$-module, we define $\operatorname{disc}(\mathcal{O})\in\mathbf{Z}$ as the discriminant of a $\mathbf{Z}$-basis for $\mathcal{O}$. In particular, $\operatorname{disc}(\mathcal{O}_{K})$ is known as the discriminant of the number field $K$ and it is an important arithmetic invariant encoding information about $K$. If $\alpha\in\mathcal{O}$, then we have
@@
The index $[\mathcal{O}:\mathbf{Z}[\alpha]]$ can be determined from an explicit basis for $\mathcal{O}$ in terms of $\alpha$. Knowing the discriminant of the power basis $1,\alpha,\ldots,\alpha^{n-1}$ then allows us to compute $\operatorname{disc(\mathcal{O})}$.
@@
In the remainder of this section, let $R$ be an integral domain, $f,g$ be polynomials of degree $n$ and $m$ respectively with coefficients given by $f(X)=a_{n}X^{n}+\cdots+a_{1}X+a_{0}\in R[X]$ and $g(X)=b_{m}X^{m}+\cdots+b_{1}X+b_{0}\in R[X]$. Our path towards formalized computation of the discriminant starts with the resultant $\operatorname{Res}(f,g)$ defined as the determinant of the Sylvester matrix of dimension $(m+n)\times(m+n)$, determined by the coefficients of $f$ and $g$ (see (Lang, 2002, IV, §8) for a transposed version with $f$ and $g$ interchanged and coefficients reversed); we work out an example in Appendix B. Let $f$ be a polynomial with formal derivative $f^{\prime}$, then $\operatorname{Res}(f,f^{\prime})$ is divisible by the leading coefficient $a_{n}$ of $f$; one sees this by inspection of the last row of the Sylvester matrix, which contains two nonzero entries: $a_{n}$ and $na_{n}$. We define the discriminant of a polynomial $f$ with leading coefficient $a_{n}$ such that $a_{n}\operatorname{Disc}(f)=(-1)^{n(n-1)/2}\operatorname{Res}(f,f^{\prime})$, by modifying the Sylvester matrix to divide the last row by $a_{n}$, and multiplying the determinant of this modified matrix by the desired sign.
@@
$\operatorname{Res}(f,g)=a_{n}^{m}b_{m}^{n}\prod_{i}\prod_{j}(\alpha_{i}-\beta_{j}).$
Note that the assumption of Theorem 6.1 always applies after mapping $f$ and $g$ to the algebraic closure of the fraction field of $R$, and this map will preserve the resultant, so we will assume that $f$ and $g$ split completely. It then follows from the product rule and elementary algebraic manipulations that, for $f$ monic with roots $\alpha_{i}$, $\prod_{i}\prod_{j>i}(\alpha_{j}-\alpha_{i})^{2}=\operatorname{Disc}(f)$. Since the former term is shown in Mathlib to equal the discriminant of the power basis $1,\ldots,\alpha^{n-1}$, and the latter term is computable, this gives us a way to compute, given a basis for $\mathcal{O}_{K}$ and using (5), the discriminant of a number field.
@@
Next, we may take $f$ and $g$ monic and view $\operatorname{Res}(f,g)$ and $\prod_{i}\prod_{j}(\alpha_{i}-\beta_{j})$ as polynomials in $R[\alpha_{1},\dots,\alpha_{n},\beta_{1},\dots,\beta_{m}]$. We showed that these are indeed homogeneous polynomials, of degree $mn$, where an essential ingredient was characterizing the homogeneous polynomials $p\in R[x,y,\dots,z]$ as those polynomials where $p(cx,cy,\dots,cz)=c^{k}p(x,y,\dots,z)$, assuming $R$ is infinite. The coefficient of $(\alpha_{1}\alpha_{2}\cdots\alpha_{n})^{m}$ in $\operatorname{Res}(f,g)$ and in $\prod_{i}\prod_{j}(\alpha_{i}-\beta_{j})$ equals $1$, so if the two expressions are equal up to multiplication by $c$, we must have $c=1$ and the expressions are actually equal.
@@
Although this approach does not differ significantly from one found in textbooks such as Lang’s Algebra (Lang, 2002), a particular source of difficulty we had to deal with was the textbooks’ implicitly shifting point of view between different polynomial rings. For example, Lang argues that $\operatorname{Res}(f,g)$ is homogeneous of degree $m$ as a polynomial in $R[a_{0},\dots,a_{n}]$ and of degree $n$ as a polynomial in $R[b_{0},\dots,b_{m}]$, but that $\alpha_{i}-\beta_{j}$ divides $\operatorname{Res}(f,g)$ in $R[\alpha_{1},\dots,\alpha_{n},\beta_{1},\dots,\beta_{m}]$. Rather than having to repeat the boilerplate involved with a change of rings repeatedly throughout our proofs, we chose to adapt the proofs from the literature to remain within one ring. Our formalization works exclusively in the polynomials
$R[\alpha_{1},\dots,\alpha_{n},\beta_{1},\dots,\beta_{m}]$. For example, polynomials that are expressed in terms of $a_{0},\dots,a_{n},b_{0},\dots,b_{m}$ can be translated to $R[\alpha_{1},\dots,\alpha_{n},\beta_{1},\dots,\beta_{m}]$ using the monadic bind operator (Commelin and Lewis, 2021), since each coefficient $a_{i}$ can be given as a polynomial in the roots $\alpha_{j}$.
To evaluate discriminants, we use the Polynomial.ofList definition of Section 4.1. Expressions such as resultant (ofList [5, 4, 3, 2, 1]) (ofList [4, 6, 6, 4]) are computable, although we run into a subsequent error: the implementation of determinants in Mathlib is not optimized for computation and therefore causes a stack overflow. However, we were still able to compute the discriminant of degree $3$ polynomials such as $f=X^{3}-30X-80$ and, after formally verifying an integral basis for the ring of integers of $K=\mathbf{Q}[X]/\langle f\rangle$, proved
@@ 7. Formally Verifiying LMFDB Entries
A popular publicly available online database containing number theoretic data is the L-functions and modular forms database (LMFDB) (LMFDB Collaboration, 2024). It gathers data from various sources, relying on several programming languages and CASes. Much of the data is currently beyond the reach of formalization, but our work makes it possible to formally verify some nontrivial entries for number fields of arbitrary degree. In particular, we focus on checking the integral basis for a choice of number fields up to degree $8$. To a lesser extent, we also compute the corresponding discriminant, but only up to degree $3$, as our computational methods are currently still under development. At this point the aim is not to formally verify integral bases and discriminants for a large fraction of the LMFDB, but to show the feasibility, and current computational limits, of the approach.
Concretely, there are $142$ number fields in the LMFDB which are of degree $5$, unramified outside $2,3,5$, and have (to focus on complications) non-monogenic ring of integers. (This list, up to isomorphism, is complete, but this is not our concern here.) For all these $142$ number fields we successfully used our tools to formally verify the given integral basis. The most time-consuming step when checking these proofs appears to be the exponentiation step in the certificates of Section 5.6. We recall that this is done with repeated squaring and multiplication. Using a virtual machine running an AMD EPYC 7B13 processor (2.45 GHz) and 16GB of RAM, each ring of integers took, on average, around 33 seconds to check. Approximately 14% of that time is spent on the irreducibility proof for the defining polynomial. Replacing decide and decide! with native_decide whenever possible reduces the overall checking time by about 20%. During development, we encountered a couple of examples that required more than the default number of maximum heartbeats (200,000) for type-checking, and the proofs had to be manually adjusted. This was resolved after a Mathlib update.
As our discriminant computations cannot yet handle large degrees, we also tested degree $3$ cases. There are $7$ number fields in the LMFDB which are of degree $3$, unramified outside $2,3,5$, and have non-monogenic ring of integers. For all these number fields we successfully formally verify the given integral basis as well as the discriminant. We note that in all of these cases, the longer form of the certificate for $p$-maximality (given in detail in Appendix A) had to be used. In contrast, for all of degree 5 examples except one, the shorter certificate with a witness was sufficient. In this case, each ring of integers took, on average, around 28 seconds to check. Using native_decide decreases this time by around 10%.
We also successfully verified a degree 6 example and a simpler degree 8 example. Both took around 48 seconds to check. For these, native_decide reduces the checking time by 61% and 32%, respectively.
@@ 8.1. Future Work
The discriminant computations should be improved. At this point, these are directly computed from the corresponding resultant. For repeated use of discriminant evaluation of a fixed degree, general formulae for the discriminant of a degree $n$ polynomial should be formalized, up to some ‘reasonable’ $n$. The computations for the ring of integers are in a more advanced stage, though further optimization is needed to deal with higher degrees in order to be able to cover all number fields from the LMFDB. We note that much of the mathematical foundation is formalized in enough generality to allow for possible extension to more general contexts. Although not essential for database verification – since all necessary Lean files can be generated in advance– integrating SageMath and Lean, for instance by calling our SageMath scripts directly from Lean, could be convenient for on-demand computations. Finally, certifying other fundamental invariants, such as class groups and unit groups, would be very interesting. While these would constitute completely new projects, they could build on the formalized rings of integers.
@@ 8.2. Related Work
As mentioned above, some rings of integers for quadratic number fields have been formalized before (Baanen et al., 2023); this concerns both the determination as well as concrete computation. Specific instances, such as the Gaussian integers $\mathbf{Z}[i]$, have been considered in various systems; see e.g. (Eberl, 2020; Futa et al., 2012). Mathematical certificates outside of algebraic number theory have received much attention. Some examples relevant for (general) number theory and computer algebra include (Thiemann et al., 2020; Mahboubi and Sibut-Pinote, 2021). Previous work bridging Lean with Computer Algebra Systems includes the tactics polyrith, written by Dhruv Bhatia, and sageify (Baanen et al., 2023), both of which integrate with SageMath. Additionally,  (Lewis, 2017; Lewis and Wu, 2022) develop a connection between Lean and Mathematica. The use of a proof assistant as a CAS itself has also been explored in prior work. In (Dénès et al., 2012), efficient algorithms in linear algebra and polynomial arithmetic are implemented and verified in Coq. Similarly, (Aransay and Divasón, 2017, 2016) implements various linear algebra algorithms in Isabelle/HOL.
@@ 8.3. Conclusions
In total, for our developments, we wrote about 13000 lines of Lean code. Our SageMath scripts automatically generated about 32000 lines of Lean code (excluding the alternative code containing native_decide) to perform the certifications we considered.
With these kind of projects, it is very hard to reliably measure the De Bruijn factor.
Appendix A Certificate for $p$-maximality
The following certificate is guaranteed to always exist if $K$ is a number field, $\mathcal{O}^{\prime}=\mathcal{O}_{K}$, and $\mathcal{O}$ is $p$-maximal.
@@ Appendix A Certificate for $p$-maximality
Consider $\mathcal{O}$ and $\mathcal{O}^{\prime}$ subalgebras of a $\mathbf{Z}$-torsion-free commutative ring such that $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, and both of rank $r$. Let =$\mathcal{B}=\{b_{1},\ldots,b_{r}\}$ be a $\mathbf{Z}$-basis for $\mathcal{O}$. Let $m,n$ and $t$ be non-negative integers with $r=m+n$ and $r\leq p^{t}$. A certificate for $p$-maximality consists of the following data:
@@
(p : ℕ) [hpI : Fact \$ Nat.Prime p] ($\mathcal{O}$ : Subalgebra K) ($\mathcal{O}$ : Subalgebra K) (hm : $\mathcal{O}$ $\mathcal{O}$’) where
Appendix B Sylvester matrix computation examples
Consider the polynomial $f=2X^{3}-X^{2}-2X+1\in\mathbf{Z}[X]$, which has derivative $f^{\prime}=6X^{2}-2X-2$. The Sylvester matrix for $f$ and $f^{\prime}$ is given by:
$A$ $=\begin{pmatrix}-2&0&0&1&0\\ -2&-2&0&-2&1\\ 6&-2&-2&-1&-2\\ 0&6&-2&2&-1\\ 0&0&6&0&2\end{pmatrix}.$
The resultant is defined as the determinant of the Sylvester matrix, giving us $\operatorname{Res}(f,f^{\prime})=|A|=-72$.
To compute the discriminant of $f$ we modify the Sylvester matrix by dividing out the leading coefficient of $f$, $2$, in the last row:
$A$ $=\begin{pmatrix}-2&0&0&1&0\\ -2&-2&0&-2&1\\ 6&-2&-2&-1&-2\\ 0&6&-2&2&-1\\ 0&0&3&0&1\end{pmatrix}.$
We obtain $\operatorname{Disc}(f)=(-1)^{\deg f(\deg f-1)/2}|A^{\prime}|=(-1)^{3}\cdot-36=36$.
Passing to the field of fractions, we can divide $f$ by its leading coefficient to get the monic polynomial $g=X^{3}-\frac{1}{2}X^{2}-X+\frac{1}{2}\in\mathbf{Q}[X]$, with discriminant computed as before $\operatorname{Disc}(g)=\frac{9}{4}=\left(\frac{1}{2}\right)^{2(\deg f-1)}\operatorname{Disc}(f)$. The roots of $f$ and $g$ in $\mathbf{Q}$ are given by $X\in\{-1,\frac{1}{2},1\}$, and we find as expected that the product of root differences equals the discriminant of $g$:
$\left(\frac{1}{2}--1\right)^{2}\left(1--1\right)^{2}\left(1-\frac{1}{2}\right)^{2}=\frac{9}{4}.$
Data-Availability Statement
Full source code of our formalization and SageMath scripts are available online at https://github.com/alainchmt/RingOfIntegersProject and are archived as (Chavarri Villarello et al., 2024).
Acknowledgments
Alain Chavarri Villarello was funded by NWO Vidi grant No. 613.009.143, Formalizing Diophantine algorithms. We thank Assia Mahboubi, Filippo A. E. Nuccio and the anonymous referees for their helpful comments.